OpenStudy (anonymous):

WHO IS GREAT AT 8TH GRADE MATH!? I WILL GIVE METALS!

OpenStudy (anonymous):

depends on what the question is

OpenStudy (anonymous):

ok

OpenStudy (holly00d1248):

Me wat is it

OpenStudy (studygurl14):

What's your question @Mlink1290 ? Just post it.

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

I have 5 questions

OpenStudy (anonymous):

A system of equations is given below: y = –2x + 1 6x + 2y = 22 Which of the following steps could be used to solve by substitution? (4 points) –2x + 1 = 6x + 2y 6x + 2(–2x + 1) = 22 6(–2x + 1) + 2y = 22 6(y = –2x + 1)

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

ok beutiful

OpenStudy (anonymous):

first one is 24x ok

OpenStudy (anonymous):

6x + 2y = 10

OpenStudy (studygurl14):

If you have y = -2x + 1 and 6x + 2y = 22 you can plug in -2x + 1 for y in the second equation (substitution). So what's the answer @Mlink1290 ?

OpenStudy (anonymous):

@Daniel13519 thats not a option

OpenStudy (anonymous):

hey mink is mine @mink1290

OpenStudy (anonymous):

no its not an option either

OpenStudy (anonymous):

heres a screen shot of the question

OpenStudy (anonymous):

Theres the question..

OpenStudy (studygurl14):

Did you read my answer @Mlink1290 ?

OpenStudy (anonymous):

yes I dont understand and my internet crashed lol sorry

OpenStudy (studygurl14):

If y equals -2x + 1 and you have the equation 6x + 2y = 22, you can substitute -2x + 1 for y in the equation. y = -2x + 1 6x + 2y = 22 6x + 2(-2x + 1) = 22 Like that.